# InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection

### v5

@@ B.1 Hyperparameter details
Table 6: Hyperparameter settings.
@@
Figure 6: Hyperparameter study curves for w/ T2V. Panel (d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.
B.2 Hyperparameter study
@@
B.3 Empirical study of self-attention fine-tuning and interaction modes
@@
Table 7: Average results of fine-tuning different weight matrices $\mathbf{W}$ across four baseline methods.
@@
Table 8: Average results of four baseline methods for fine-tuning different weight matrices $\mathbf{W}$.
Appendix C More Visual Examples
Figure 7 presents additional examples illustrating the focus differences between InterCLIP and CLIP. These visualizations highlight InterCLIP’s improved ability to capture sarcasm-related cues by focusing on relevant areas in the images.
Figure 7: Additional visual examples showcasing InterCLIP’s improved focus on sarcasm-related visual cues compared to CLIP.
Appendix D Extended Experiments
To further verify the performance of our framework, we conducted an extended set of experiments.
@@ Appendix D Extended Experiments
Table 9: Statistics of DocMSU.
@@
Table 10: Results of the extended experiments. Underline results denote the compared SOTA baseline, boldface highlights results that surpass the baseline, and * indicates results sourced from Du et al. (2024).
D.1 Benchmark
DocMSU Du et al. (2024) is a recently introduced multi-modal sarcasm benchmark designed specifically for long-text analysis. It facilitates the evaluation of multi-modal sarcasm comprehension as well as detection tasks. In this study, we concentrate on the multi-modal sarcasm detection task. The benchmark statistics can be found in Table 9.
D.2 Baselines
We follow the evaluation protocol of Du et al. (2024). We compare against unimodal baselines: BERT-base (text-only) Devlin (2018) and Swin Transformer (image-only) Liu et al. (2021). For multi-modal approaches, we include CLIP Radford et al. (2021), Vision-and-Language Transformer (ViLT) Kim et al. (2021), and CMGCN Liang et al. (2022). The method proposed by Du et al. (2024) is taken as the state-of-the-art baseline.
D.3 Results
@@
Our InterCLIP-MEP framework demonstrates strong performance across various configurations, as shown in Table 10. In particular, the w/o Inter and w/ V2T variants consistently achieve higher F1 scores compared to the baselines, thereby showcasing their robustness in handling multi-modal sarcasm detection tasks. Notably, all variants achieve nearly perfect recall, highlighting their outstanding capability in accurately identifying sarcasm across diverse datasets. The w/o Inter variant also achieves the highest accuracy, further demonstrating its effectiveness and precision.
Overall, the comprehensive results in Table 10 affirm the unparalleled effectiveness of InterCLIP’s modality interaction mechanism and our proposed memory-enhanced predictor (MEP), particularly in surpassing existing state-of-the-art methods in both accuracy and F1 score, thus setting a new benchmark in the field.
Appendix E Other Related Works
E.1 CLIP adaptation
The Contrastive Language-Image Pretraining (CLIP) model Radford et al. (2021) excels in vision-language tasks. Adapting CLIP for specific domains has shown substantial improvements, as demonstrated by Li et al. (2022) for phrase localization, Liang et al. (2023) for open-vocabulary semantic segmentation, and Wang et al. (2023) for action recognition. In this work, inspired by Ganz et al. (2024), we conditionally enhance both the text and vision encoders of CLIP, making it more effective in capturing the interplay between text and images to identify multi-modal sarcasm cues. Unlike Ganz et al. (2024), who focused solely on embedding text representations into the vision encoder, we also explore embedding image representations into the text encoder. Furthermore, their approach is limited to general classification tasks and does not address the complexities of multi-modal sarcasm detection.
E.2 Memory-enhanced prediction
Inspired by cognitive science Stokes (2015); Baddeley (2000), memory has been introduced to enhance neural networks Weston et al. (2014); Sukhbaatar et al. (2015). Several studies Wu et al. (2018); Wen et al. (2023) have used memory mechanisms to improve model training, and some Zhang et al. (2024); Wei et al. (2024) leverage memory to store historical knowledge, enhancing prediction accuracy. In this work, we introduce a memory-enhanced predictor for multi-modal sarcasm detection. In contrast to other methods, our memory dynamically updates during testing, utilizing relevant historical information for improved accuracy and robustness.
Appendix F List of Symbols
In Table 11, we have listed the main symbols used in the paper and their descriptions.
Symbol Description
$T$ $T$ denotes a short text.
$I$ $I$ represents an image.
$\mathcal{P}$ $\mathcal{P}$ denotes a text-image pair $(T,I)$.
$\mathcal{T}$ $\mathcal{T}$ denotes CLIP’s text encoder.
$\mathcal{V}$ $\mathcal{V}$ denotes CLIP’s vision encoder.
Table 11: List of symbols

### v7

@@ 4.2. Implementation Details
4.3. Hyperparameter Details
We summarize the hyperparameters involved in InterCLIP-MEP and their descriptions in Table 4. The hyperparameter settings for obtaining the main results in the paper are summarized in Table 4. For each experiment, we condition only the top four layers of the self-attention modules, with the projection dimension $d_{f}$ set to 1024. We set the LoRA rank $r$ to 8, fine-tuning the self-attention weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$. For the memory size $L$, we select the optimal size from a fixed set of candidate values $\mathbf{L}$. For other baseline methods, we follow the optimal hyperparameter settings they reported.
4.4. Main Results
To begin our experimental analysis, we first evaluate the overall effectiveness of the proposed InterCLIP-MEP framework. To further assess its robustness, we design a control variant, denoted as w/o Inter, where the original CLIP is used as the backbone in place of InterCLIP. This variant is compared against three interaction modes of InterCLIP: w/ V2T, w/ T2V, and w/ TW. The overall results and baseline comparisons are summarized in Table 5, while the effectiveness of different interaction modes is further analyzed in Table 6.
Performance on MMSD2.0
As shown in Table 5, our framework consistently surpasses or matches state-of-the-art baselines on MMSD2.0, confirming the effectiveness of the proposed training strategy and the Memory-Enhanced Predictor (MEP). Table 6 further compares different interaction settings. Both w/ V2T and w/ T2V outperform the w/o Inter variant, indicating that InterCLIP more effectively captures cross-modal sarcasm cues than the original CLIP backbone. Among them, w/ T2V yields the best results. This is because projecting compact, low-dimensional text embeddings into the higher-dimensional vision space allows redundancy in visual features to accommodate textual semantics, preserving information more effectively. In contrast, projecting rich, high-dimensional vision embeddings into the text space requires compression into a compact semantic space, which inevitably loses details and hampers alignment. This asymmetry explains why T2V consistently outperforms V2T. Meanwhile, the w/ TW configuration performs worse than either single-direction variant, as simultaneously embedding both modalities exacerbates the alignment asymmetry and further increases the learning burden. In summary, InterCLIP with T2V interaction, together with our lightweight training strategy and MEP, achieves the most promising results on MMSD2.0, underscoring the robustness and adaptability of our framework in modeling subtle text–image interactions.
Table 5. Baseline comparison on MMSD2.0 and MMSD, where InterCLIP-MEP uses the best interaction mode for each dataset.
Method MMSD2.0 MMSD
Acc. (%) F1 (%) P (%) R (%) Acc. (%) F1 (%) P (%) R (%)
Text
TextCNN (kim-2014-convolutional) 71.61* 69.52* 64.62* 75.22* 80.03* 75.32* 74.29* 76.39*
Bi-LSTM (zhou-etal-2016-attention) 72.48* 68.05* 68.02* 68.08* 81.90* 77.53* 76.66* 78.42*
SMSD (xiong2019sarcasm) 73.56* 69.97* 68.45* 71.55* 80.90* 75.82* 76.46* 75.18*
RoBERTa (liu2019roberta) 79.66* 76.21* 76.74* 75.70* 93.97* 92.45* 90.39* 94.59*
Image
ResNet (he2015deep) 65.50* 57.58* 61.17* 54.39* 64.76* 61.53* 54.41* 70.80*
ViT (dosovitskiy2020image) 72.02* 69.72* 65.26* 74.83* 67.83* 63.40* 57.93* 70.07*
Text-Image
HFM (Cai2019MultiModalSD) 70.57* 66.88* 64.84* 69.05* 83.44* 80.18* 76.57* 84.15*
Att-BERT (pan-etal-2020-modeling) 80.03* 77.04* 76.28* 77.82* 86.05* 82.92* 80.87* 85.08*
CMGCN (Liang2022MultiModalSD) 79.83* 76.90* 75.82* 78.01* 86.54* 84.09* -* -*
HKE (Liu2022TowardsMS) 76.50* 72.25* 73.48* 71.07* 87.36* 72.25* 81.84* 86.48*
DIP (wen2023dip) 80.59 78.23 75.52 81.14 89.59 87.17 87.76 86.58
DynRT (Tian2023DynamicRT) 70.37 68.55 63.02 75.15 93.59 91.93 90.30 93.62
Multi-view CLIP (qin2023mmsd2) 85.64* 84.10* 80.33* 88.24* 88.33* 85.55* 82.66* 88.65*
G2SAM (wei2024g2sam) 79.43 78.07 72.04 85.20 90.48 88.48 87.95 89.02
InterCLIP-MEP (Ours) 86.72 85.61 80.20 91.80 93.94 92.54 90.77 94.37
    We use * to indicate that the results are taken from (qin2023mmsd2). indicates that results are not reported. $\dagger$ indicates our reproduced results. Underlined values represent the best multi-modal baseline for comparison. Bold values indicate those that surpass the underlined baseline. The InterCLIP-MEP results use the T2V interaction on MMSD2.0 and the w/o Inter setting on MMSD. These choices follow our analysis (Section 4.4, Table 6), which shows T2V is most effective on MMSD2.0, while MMSD’s spurious textual cues make w/o Inter more suitable.
Table 6. Performance comparison of InterCLIP-MEP under different interaction settings.
Method MMSD2.0 MMSD
Acc. (%) F1 (%) P (%) R (%) Acc. (%) F1 (%) P (%) R (%)
InterCLIP-MEP
InterCLIP-MEP w/ RoBERTa
    Bold values indicate those that surpass the underlined baseline. $L_{2}$ for MMSD2.0 and $L_{1}$ for MMSD denote the optimal MEP memory sizes.
Performance on MMSD
Table 7. Performance comparison on the MMSD dataset in terms of Macro-average F1 score, following the standard evaluation protocol used in recent literature (zhuang2025multi). Results marked with * are reported from (zhuang2025multi).
Method Macro-F1 (%)
DIP (wen2023dip)* 89.01
G2SAM (wei2024g2sam)* 89.65
KFGC-Net (zhuang2025multi)* 91.02
InterCLIP-MEP 93.72
As shown in Table 5, the RoBERTa-based text baseline significantly outperforms other methods on MMSD due to spurious textual cues, which allow models to achieve high accuracy using text features alone (qin2023mmsd2). Consequently, multi-modal models such as DynRT (Tian2023DynamicRT), G2SAM (wei2024g2sam), and DIP (wen2023dip), which rely on RoBERTa or BERT for text encoding, also obtain strong results on MMSD but exhibit a substantial performance drop on MMSD2.0. To further investigate this effect, we introduce an additional variant, InterCLIP-MEP w/ RoBERTa, by replacing the original text encoder with RoBERTa (Table 6). Although this variant achieves state-of-the-art performance on MMSD, it disrupts InterCLIP’s modality alignment and thus leads to a noticeable decline on MMSD2.0. These findings suggest that MMSD’s text data contains spurious cues that encourage over-reliance on textual signals, whereas MMSD2.0, being cleaned, requires stronger multi-modal reasoning. We also observe that the w/ V2T variant performs best in both InterCLIP-MEP and InterCLIP-MEP w/ RoBERTa, highlighting the model’s tendency to over-depend on textual features. Following (zhuang2025multi), we also report the Macro-F1 score on the MMSD dataset to ensure consistency with community evaluation standards. As shown in Table 7, InterCLIP-MEP achieves the highest Macro-F1 of 93.72%, outperforming DIP (wen2023dip), G2SAM (wei2024g2sam), and KFGC-Net (zhuang2025multi).
Table 8. Efficiency comparison of different methods.
Multi-view CLIP  85.64  165  488  51  15.59
DIP  80.59  196  OOM  OOM  OOM
G2SAM  79.43  116  90  13  18.32
DynRT  70.37  25  370  26  8.03
InterCLIP-MEP  86.72  8  55  6  6.14
    To demonstrate the efficiency of InterCLIP-MEP, we selected several recent baselines for comparison. The analysis was conducted using the MMSD2.0 dataset on a single NVIDIA RTX 4090 GPU with a batch size of 128. In the table, Fitting Time / Epoch indicates the time required for each epoch during training and validation and OOM indicates Out of Memory, referring to GPU memory overflow.
Efficiency Comparison.
Our training strategy demonstrates both remarkable effectiveness and outstanding efficiency. To validate this, we conducted a comprehensive comparative analysis against leading state-of-the-art methods, as detailed in Table 8. For instance, the Multi-view CLIP method (qin2023mmsd2) employs a multi-layer Transformer encoder for feature fusion, which, while effective, introduces a significant number of trainable parameters. This results in slower training and inference speeds and greater memory consumption. Similarly, the DIP method (wen2023dip) caches historical samples during training, which hinders its ability to support large-batch training under limited resource conditions. In contrast, our method operates with a batch size of 128 while utilizing a significantly smaller number of trainable parameters, which translates to notably faster training and validation cycles. Furthermore, by incorporating minimal parameter modifications to adapt CLIP and utilizing simple yet effective linear layers for representation fusion, our approach achieves superior inference speeds and drastically reduced memory consumption. These results highlight the practicality of our framework, establishing it as a benchmark for both computational efficiency and performance in multi-modal sarcasm detection.
Table 9. Ablation study of InterCLIP-MEP.
w/o Inter w/ TW w/ V2T w/ T2V
Variant Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%)
Baseline 86.05 84.81 85.51 84.26 86.26 85.00 86.72 85.61
w/o Proj 85.76 84.43 85.43 84.05 85.68 84.22 86.22 84.51
w/o MEP 85.39 83.99 85.22 83.79 86.26 84.78 86.26 84.82
w/o LoRA 82.44 77.73 76.42 74.37 73.31 72.22 75.13 71.79
    Baseline denotes the results without ablation.
4.5. Analysis of InterCLIP-MEP
Ablation study.
We remove the projection module $\mathcal{F}_{p}$ and train only the classification module $\mathcal{F}_{c}$ for prediction, denoted as w/o Proj. To test the necessity of using LoRA (edward2022lora) for fine-tuning, we keep the rest of InterCLIP-MEP unchanged and freeze all self-attention weights of InterCLIP, denoted as w/o LoRA, always selecting the optimal memory size $L$ for the MEP during inference. To evaluate the effectiveness of the MEP, we train both $\mathcal{F}_{p}$ and $\mathcal{F}_{c}$ but use only $\mathcal{F}_{c}$ during inference, denoted as w/o MEP. Table 9 reports all results. All variants show performance declines compared to the baseline, demonstrating the importance of each module in InterCLIP-MEP. For w/ TW and w/o Inter, the w/o MEP variant performed worse than the w/o Proj variant. However, for w/ T2V and w/ V2T, the w/o MEP variant performs better than the w/o Proj variant. This suggests that backbones with strong image-text interaction capabilities benefit from training the classification module $\mathcal{F}_{c}$ along with the projection module $\mathcal{F}_{p}$, even without using MEP during inference. We also find that not using LoRA to fine-tune the self-attention modules results in significant performance loss, indicating that the original CLIP’s vanilla space is not suitable for the sarcasm detection task.
@@ 4.5. Analysis of InterCLIP-MEP
Figure 4. Hyperparameter study curves for w/ T2V. Panel (d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.
Figure 5. Case study of InterCLIP-MEP. In the figure, the emojis in the Sample column denote the ground-truth labels from the dataset. MEP represents the labels predicted by the memory-enhanced predictor, and $\mathcal{F}_{c}$ represents the labels predicted by the classification module.
Hyperparameter study.
@@
Figure 6. Visual examples showcasing InterCLIP’s improved focus on sarcasm-related visual cues compared to CLIP. In the figure, the first column shows the original input images, the second and third columns present the GradCAM visualizations from InterCLIP and the original CLIP respectively, and the last column provides human-readable explanations comparing their behaviors.
Case study.
As shown in Figure 5, we provide illustrative examples to examine both the strengths and the limitations of InterCLIP-MEP. In the first two cases, direct predictions from the classification module $\mathcal{F}_{c}$ yield uncertain and incorrect outcomes with relatively high entropy, whereas the Memory-Enhanced Predictor (MEP) leverages historical sample knowledge to successfully correct the results. These examples highlight the robustness and effectiveness of our design.
At the same time, we also include two challenging failure cases. In the third example, $\mathcal{F}_{c}$ makes a correct prediction, while MEP mistakenly overrides it. This situation arises when the memory stores similar but misleading representations, which can bias the final decision. In the fourth example, both $\mathcal{F}_{c}$ and MEP fail, reflecting the intrinsic difficulty of multi-modal sarcasm detection when subtle ironic cues are either extremely implicit or context-dependent beyond the available modalities. These observations naturally raise two questions: how does the memory mechanism handle potential error propagation, and how robust is it under distribution shifts? To answer these, we provide dedicated analyses in Section 5.3 and Section 5.4, where we examine these challenges in a systematic manner.
With these results in mind, it is important to note that the identified failures do not undermine the overall effectiveness of InterCLIP-MEP, but rather point to promising directions for refinement. For instance, incorporating context-aware memory update strategies or external commonsense knowledge may further reduce such errors. Overall, the inclusion of both successful and unsuccessful cases provides a more balanced perspective, demonstrating that while our method substantially improves robustness, multi-modal sarcasm detection remains a highly challenging task.
Visualization.
To further validate that InterCLIP is capable of more effectively capturing the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we use GradCAM (Selvaraju2017GradCAM) to visualize the areas of focus during the inference process of the visual model in Figure 6. These visualizations highlight InterCLIP’s improved ability to capture sarcasm-related cues by focusing on relevant areas in the images.
4.6. Empirical Study of Self-attention Fine-tuning and Interaction Modes
@@
Table 10. Average results of fine-tuning different weight matrices $\mathbf{W}$ across four baseline methods.
@@
Table 11. Average results of four baseline methods for fine-tuning different weight matrices $\mathbf{W}$.
@@
Table 12. Statistics of DocMSU.
@@
5. Extended Experiments
To further verify the performance and robustness of our framework, we extended our evaluation in three complementary directions. First, we conducted experiments on a document-level benchmark to examine its capability in handling long-text sarcasm detection. Second, we performed low-resource experiments to assess generalization and adaptability under data scarcity. Third, we systematically investigated the impact of our memory mechanism under error propagation and distribution shift conditions, thereby providing deeper insights into its stability and reliability.
5.1. Document-level Benchmark
DocMSU Benchmark
DocMSU (du2024docmsu) is a recently introduced multi-modal sarcasm benchmark designed specifically for long-text analysis. It facilitates the evaluation of multi-modal sarcasm comprehension as well as detection tasks. In this study, we concentrate on the multi-modal sarcasm detection task. The benchmark statistics can be found in Table 12.
Baselines
We follow the evaluation protocol of (du2024docmsu). We compare against unimodal baselines: BERT-base (text-only) (devlin2019bert) and Swin Transformer (image-only) (liu2021swin). For multi-modal approaches, we include CLIP (radford2021learning), Vision-and-Language Transformer (ViLT) (kim2021vilt), and CMGCN (Liang2022MultiModalSD). The method proposed by (du2024docmsu) is taken as the state-of-the-art baseline.
Results
Table 13. Results on DocMSU.
@@ Results
    Underline results denote the compared SOTA baseline, boldface highlights results that surpass the baseline, and * indicates results sourced from (du2024docmsu).
@@
Our InterCLIP-MEP framework demonstrates strong performance across various configurations, as shown in Table 13. In particular, the w/o Inter and w/ V2T variants consistently achieve higher F1 scores compared to the baselines, thereby showcasing their robustness in handling multi-modal sarcasm detection tasks. Notably, all variants achieve nearly perfect recall, highlighting their outstanding capability in accurately identifying sarcasm across diverse datasets. The w/o Inter variant also achieves the highest accuracy, further demonstrating its effectiveness and precision. Overall, the comprehensive results in Table 13 affirm the unparalleled effectiveness of InterCLIP’s modality interaction mechanism and our proposed memory-enhanced predictor (MEP), particularly in surpassing existing state-of-the-art methods in both accuracy and F1 score, thus setting a new benchmark in the field.
5.2. Low-resource Setting
To assess the robustness of InterCLIP-MEP under low-resource conditions, we followed the experimental protocol of Multi-view CLIP (qin2023mmsd2) and conducted experiments on the MMSD2.0 dataset. Specifically, we subsampled the training set by selecting both positive and negative samples at equal intervals of 10%, and employed the hyperparameter configuration that yielded the best performance in Table 5. The models were then evaluated on the test set, and the results are presented in Figure 9.
Figure 7. Accuracy on MMSD2.0 with different training ratios.
Figure 8. Accuracy under different levels of pseudo-label noise. The dashed line indicates the result without MEP.
Figure 9. Evaluation on a mixed MMSD2.0–DocMSU test set illustrates that interactive representations and the memory mechanism jointly enhance robustness under distributional shifts.
As shown in the figure, when 50% or more of the training data is available, InterCLIP-MEP achieves a clear advantage over Multi-view CLIP. This indicates that our framework is capable of leveraging moderately limited supervision to reconstruct a more effective multi-modal representation space. Moreover, while our method performs slightly worse than Multi-view CLIP when training with less than 50% of the data, it exhibits an emergent performance gain once this threshold is reached. Specifically, the accuracy of InterCLIP-MEP increases by 14.24% from its lowest to highest point, whereas Multi-view CLIP only improves by 2.91%.
We attribute this phenomenon to the distinct training strategies. Multi-view CLIP relies on original CLIP representations, which provide relatively strong performance even with limited supervision but yield diminishing returns as more data becomes available. In contrast, InterCLIP-MEP employs lightweight fine-tuning to restructure the representation space of CLIP, making it more aligned with the specific requirements of multi-modal sarcasm detection. This process requires a sufficient amount of data to stabilize the new cross-modal alignment, but once the data threshold is met, the framework benefits substantially more from additional training samples.
5.3. Error Propagation Robustness Study
To further assess the robustness of the proposed memory mechanism against error propagation, we designed a stress test by randomly flipping a certain proportion of pseudo-labels generated by the classification module. This operation introduces deliberately corrupted low-entropy predictions into the memory, thereby simulating potential error accumulation during inference. For consistency, all hyperparameters in this experiment follow the optimal settings reported in Table 5.
The results are shown in Figure 9. Even when up to 40% of pseudo-labels are randomly flipped, the accuracy of InterCLIP-MEP remains consistently above the baseline system without memory (86.26%). The observed variations are minimal, indicating that the memory-enhanced predictor maintains stable performance under noisy pseudo-labels. These findings demonstrate that the memory-enhanced predictor does not suffer from cascading error amplification. Instead, it maintains stable performance under noisy pseudo-labels, confirming that the dynamic update and entropy-based filtering strategy effectively suppress the influence of mislabeled entries.
5.4. Robustness under Distributional Shift
To rigorously examine robustness under distributional shifts, we compared InterCLIP-MEP, its variant without the memory mechanism (w/o MEP), and the state-of-the-art Multi-view CLIP baseline. All models were trained on the MMSD2.0 (qin2023mmsd2) training set and evaluated on a mixed testing set constructed by shuffling the MMSD2.0 and DocMSU (du2024docmsu) test sets. This heterogeneous evaluation setting introduces significant domain variation beyond the training distribution and provides a stringent test of model stability.
As shown in Figure 9, InterCLIP-MEP attains the highest accuracy of 73.96%, substantially outperforming its variant without the memory mechanism (63.97%) and Multi-view CLIP (58.62%). The nearly 10% improvement over the w/o MEP variant demonstrates that the dynamic memory mechanism significantly mitigates performance degradation by suppressing error accumulation and leveraging historical knowledge during inference. Moreover, even without the memory component, InterCLIP itself still surpasses Multi-view CLIP by more than 5%, highlighting the effectiveness of the interactive representation design. These results confirm that InterCLIP-MEP achieves superior stability and adaptability when faced with distributional shifts.
6. Limitations and Future Work
While InterCLIP-MEP demonstrates strong robustness and consistent improvements across multiple evaluations, several limitations remain that highlight promising directions for future research.
Limitations.
Despite its effectiveness, InterCLIP-MEP can be sensitive to memory noise when misleading or weakly relevant representations are retrieved, which may bias the inference process and occasionally override correct predictions. Moreover, the framework still struggles with highly implicit or context-dependent sarcasm, especially when such cues rely on external commonsense or situational background that is not explicitly encoded in the available modalities. In addition, the current design assumes a relatively static memory and homogeneous data distribution, which may limit adaptability when facing unseen domains or evolving data patterns.
Future Work.
Future studies may explore context-aware memory updating strategies that dynamically adjust to semantic drift and data uncertainty. Integrating external commonsense or world knowledge could further enhance the model’s ability to capture subtle pragmatic cues and hidden irony. Another potential direction is to introduce memory regularization or selective forgetting mechanisms to mitigate the effect of noisy retrievals and reduce error propagation under domain shifts. Finally, extending the proposed memory-augmented paradigm to other multimodal reasoning tasks, such as emotion understanding or visual entailment, could provide a broader understanding of how long-term memory interacts with perception and language in open-world scenarios.
7. Conclusion
This paper introduces InterCLIP-MEP, an innovative framework designed to address the critical challenges of modeling subtle text-image interactions and mitigating prediction uncertainty in multi-modal sarcasm detection. The proposed framework incorporates Interactive CLIP (InterCLIP), which integrates cross-modal information directly into text and image encoders. This enhances the model’s ability to interpret multi-modal sarcasm cues. Furthermore, a Memory-Enhanced Predictor (MEP) is developed to dynamically utilize knowledge from historical samples, resulting in a more robust and adaptive inference process. Extensive evaluations conducted on the MMSD, MMSD2.0, and DocMSU benchmarks demonstrate that InterCLIP-MEP achieves state-of-the-art performance while substantially reducing computational overhead. By requiring fewer trainable parameters and less GPU memory, InterCLIP-MEP offers a lightweight, efficient, and scalable solution, setting a new benchmark for multi-modal sarcasm detection.
Acknowledgements
The research presented in this paper was funded by the Anhui Provincial Natural Science Foundation under Grant No. 2308085MF220. It also received additional support from the Anhui University Natural Science Foundation under Grant No. 2023AH050914.